OpenStudy (anonymous):

My last question for the night, review is killing me: Simplify: (4a^3) ^2 (3a) ^2. I understand how to simplify, but on my worksheet, the second part was written smaller...

OpenStudy (anonymous):

\[144a ^{8}\]

OpenStudy (anonymous):

Thank you! Would you mind explaining how you got that? I understand part of it.

OpenStudy (anonymous):

4^2 = 16 (a^3)^2 = a^6 3 ^ 2 = 9 (a)^2 = a ^ 2 16 * 9 = 144 a^6 * a ^ 2 = a^8 144a^8

OpenStudy (anonymous):

Thank you so much! I know it's a rather simple simplifying equation, but it's harder to remember these when you haven't done them for three months!